# Edge-weighted Online Stochastic Matching Under Jaillet-Lu LP

### v1

@@
The online stochastic matching problem was introduced by Feldman et al. [7], together with the $(1-\frac{1}{e})$-competitive Suggested Matching algorithm. In the most general edge-weighted setting, this ratio has not been improved for more than one decade, until recently Yan [21] beat the $1-\frac{1}{e}$ bound and Qiu et al. [19] further improved the ratio to $0.650$. Both of these works measure the online competitiveness against the offline LP relaxation introduced by Jaillet and Lu [14]. This LP has also played an important role in other settings since it is a natural choice for two-choices online algorithms.
In this paper, we propose an upper bound of $0.663$ and a lower bound of $0.662$ for edge-weighted online stochastic matching under Jaillet-Lu LP. First, we propose a hard instance and prove that the optimal online algorithm for this instance only has a competitive ratio $<0.663$. Then, we show that a near-optimal algorithm for this instance can be generalized to work on all instances and achieve a competitive ratio $>0.662$. It indicates that more powerful LPs are necessary if we want to further improve the ratio by $0.001$.
@@ 1 Introduction
The online bipartite matching problem was introduced by Karp, Vazirani, and Vazirani [16], motivated by applications such as online advertising. In this problem, we are given a bipartite graph. Advertisers are modeled as offline vertices on one side of the graph, which are known at the beginning. Impressions (user queries) are modeled as online vertices on the other side, which arrive one by one. The weights on the edges indicate how much the advertisers are interested in the impressions. When an online vertex arrives, the online algorithm sees its incident edges, and must immediately and irrevocably decide how to match it. Karp et al. [16] proposed the celebrated Ranking algorithm, which achieves a competitive ratio of $1-\frac{1}{e}\approx 0.632$ in the unweighted setting. This is optimal in the worst case model, where we assume the online algorithm has no prior information about the graph.
@@
However, in real-world scenarios, the algorithm often knows some prior information. For example, online advertising platforms can use historic data to predict user behaviors in the future. Feldman, Mehta, Mirrokni, and Muthukrishnan [7] introduced the online stochastic matching model, where we assume all online vertices are drawn independently from an identical distribution which is known by the algorithm in the beginning. Feldman et al. [7] proposed the Suggested Matching algorithm which is $(1-\frac{1}{e})$-competitive even in the most difficult edge-weighted setting, and then beat the $1-\frac{1}{e}$ bound in the unweighted setting under the integral assumption [^1] It means that the expected number of online vertices of each type is an integer.. Further works removed the integral assumption [18, 14, 11, 13, 20] and beat the $1-\frac{1}{e}$ bound in the vertex-weighted setting [11, 13, 20] and the free-disposal [^2] It allows us to match an offline vertex many times and only count the most valuable one in the end. edge-weighted setting [13].
@@
A common framework of the online algorithms is to solve an LP to get a fractional matching which upperbounds the offline optimal solution, and then use the fractional solution to guide the online strategy. The online competitiveness is then measured against the LP relaxation. In particular, the LP proposed by Jaillet and Lu [14] is quite useful for two-choices algorithms (which most previous algorithms are) in the sense that it limits the case that an online vertex type is mostly matched to one offline vertex in the fractional solution, forcing most online vertices to have at least two choices. Under Jaillet-Lu LP, Huang, Shu, and Yan [13] provided both a lower bound and an upper bound of the same ratio 0.7060.7060.7060.706333The exact value is $1-\frac{1}{1-\ln 2}\big{(}\frac{1}{2e}-\frac{\ln 2}{e^{2}}\big{)}$. for unweighted, vertex-weighted and free-disposal edge-weighted settings.
@@
The edge-weighted setting was known to be strictly harder than the above settings [13]. Recently, Yan [21] beat the $1-\frac{1}{e}$ bound with a competitive ratio of $0.645$ by extending the Suggested Matching algorithm to have different strategies in $3$ different time periods. Qiu, Feng, Zhou, and Wu [19] improved the ratio to $0.650$ by increasing the number of periods to $5$. Both algorithms still use the trivial strategy from Suggested Matching in more than half of the time. In contrast, instead of using several fixed strategies in different periods, the state-of-the-art Poisson Online Correlated Selection algorithm [13] for unweighted/vertex-weighted settings has a strategy which smoothly changes over time. A natural question arises: Can smoothly changed strategies help to further improve the competitive ratio in the edge-weighted setting?
@@
On the other hand, both of [21, 19] use the Jaillet-Lu LP relaxation. We know that the provably best ratio for online algorithms under Jaillet-Lu LP is $0.706$ in the other $3$ settings. What about the edge-weighted setting?
@@ 1.1 Our Results
We first propose a hard instance and show that, for this instance, the optimal online algorithm can only achieve a competitive ratio $<0.663$. Then, we extract some key properties of a near-optimal matching process of this specific instance which are sufficient to guarantee a competitive ratio $>0.662$. Finally, we propose an algorithm which maintains these properties for all instances, which is therefore $0.662$-competitive for all instances. The algorithm improves the best known ratio of $0.650$ [19], and the hardness result suggests that more powerful LPs are necessary for further improvements.
Compared with all existing algorithms for edge-weighted online stochastic matching, our algorithm has two major differences: (1) Instead of only using one or several fixed strategies, we deploy a strategy that varies from time to time; (2) When an online vertex arrives, instead of only locally looking at its neighbors, our algorithm utilizes some global information about the probability distribution of remaining offline vertices.
@@
Previous work [21] showed that, under Jaillet-Lu LP, the problem can be reduced to the case that there are only two classes of online vertex types:
@@
By carefully setting the edge weights, we show that the optimal algorithm for this instance only achieves a competitive ratio $<0.663$.
@@
We modify the optimal algorithm for the hard instance to get a near-optimal algorithm $A$, which on this instance will match each edge $e$ with probability $\Gamma x_{e}$, where $\Gamma>0.662$ and $x_{e}$ is the matching fraction of $e$ in the optimal fractional matching. We want it to be generalized to work for all instances.
First-class types are easy to deal with since they have only one possible choice when arriving, so we focus on second-class types. For each second-class type, we want to match it to each neighbor at the same rate as $A$. To make this feasible, we need an upper bound of the probability that two offline vertices have both been matched at some time $t$. We show that this probability takes its maximum value in our hard instance when we “simulate” $A$, and this finally allows us to achieve a competitive ratio of $\Gamma>0.662$.
@@ 1.2 Other Related Work
In the worst case model, the Ranking algorithm [16] and its generalization [1] achieved the optimal competitive ratio of $1-\frac{1}{e}$ for unweighted and vertex-weighted matching. For free-disposal edge-weighted matching, a sequence of works [6, 8, 3] improved the competitive ratio to $0.5239$.
In the random order model [^4] The adversary decides the graph, then online vertices arrive in a uniformly random order., the currently best ratios are $0.696$ [9, 17] for unweighted matching and $0.662$ [12, 15] for vertex-weighted matching.
In the online stochastic matching model, for unweighted and vertex-weighted matching, a sequence of works [7, 2, 18, 14, 11, 13, 20] improved the competitive ratio to $0.716$ and the upper bound to $0.823$. For free-disposal edge-weighted matching, Huang et al. [13] proposed a $0.706$-competitive algorithm. Under the integral assumption, the state-of-the-art algorithms achieved competitive ratios of $0.7299$ for unweighted matching [18, 14, 4], $0.725$ for vertex-weighted matching [14], and $0.705$ for edge-weighted matching [10, 4]. An upper bound of $0.703$ [13] for (general) edge-weighted matching inferred that it is strictly harder than the above settings. For non i.i.d. arrivals, recent works [20, 5] achieved a competitive ratio of $0.69$ for unweighted and vertex-weighted matching.
@@ 2 Preliminaries
In the Poisson arrival model, instead of assuming a deterministic number of online vertices, we assume that online vertices arrive following a Poisson process with arrival rate $\Lambda$ in the time horizon $[0,1]$. An equivalent view is that each online vertex type $i\in I$ independently follows a Poisson process with arrival rate $\lambda_{i}$. This equivalency and the continuous arrival process greatly help the analyses of online stochastic matching algorithms, after Huang and Shu [11] proved that the Poisson arrival model is asymptotically equivalent to the original model. In this paper, we present our upper and lower bounds in the Poisson arrival model. We remark that the same bounds hold in the original model with an additive term of $o(1)$ with respect to $\Lambda$. We also remark that, prior works in the original model also conventionally assume a large enough $\Lambda$, or explicitly have the $o(1)$ additive term.
@@
The (Poisson arrival version of) Jaillet-Lu LP [14] is as follows:
@@
A solution of the Jaillet-Lu LP is a fractional matching between online vertex types and offline vertices. Huang and Shu [11] proved that the optimal value of the LP, that is, the objective of the optimal fractional matching, upperbounds the expected objective of the optimal offline integral matching. In this paper, we consider the competitive ratio against the optimal value of the Jaillet-Lu LP instead of the optimal integral matching.
Without loss of generality, we assume $x_{ij}=0$ for any pair $(i,j)\notin E$. For each online vertex type $i$, define $x_{i}=\sum_{j\in J}x_{ij}$. For each offline vertex $j$, define $x_{j}=\sum_{i\in I}x_{ij}$.
@@
Yan [21] showed that, any instance $G$ can be efficiently transformed into another instance $G^{\prime}$ such that the optimal fractional matching (under Jaillet-Lu LP) of $G^{\prime}$ satisfies the following conditions:
@@
In addition, for any online algorithm which works for $G^{\prime}$, using the transformation as a preprocessing step, it works for $G$ with the same competitive ratio. So, without loss of generality, in this paper we restrict the instance to satisfy the above conditions.
@@ 3 Hardness
The optimal value of Jaillet-Lu LP on $G$ is $2\ln 2+(2-\ln 2)k$.
Now we are going to find the optimal online algorithm on $G$. Clearly, since $k\geq 1$, when a first-class online vertex arrives, we shall always match it if we can. What about a second-class online vertex? Intuitively, the decision should depend on $k$ and the current time $t$. When $t$ is small but $k$ is large, we may want to discard this online vertex even if it has unmatched neighbors, to preserve the chance of matching a second-class vertex in the future. Also notice that, the number of unmatched neighbors may also affect our decision, since it affects the chance that the offline vertex can be matched with another second-class online vertex in the future.
@@
When a second-class online vertex arrives, the algorithm can match it if $\textsc{Opt}(G^{\prime},t)+1\geq\textsc{Opt}(G,k)$, and it doesn’t matter which neighbor is matched by symmetry. When $t>t_{1}$, we know $\textsc{Opt}(G^{\prime},t)\leq 1$, then $\textsc{Opt}(G^{\prime},t)+1\geq 2\textsc{Opt}(G^{\prime},t)\geq\textsc{Opt}(G,t)$. So the algorithm can always match it when $t>t_{1}$. Then, it remains to show that $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ is decreasing in $[0,t_{1}]$. If this is true, then we can set $t_{0}$ as the boundary where $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ starts going below $1$.
@@
Note that, by Lemma 2, when $t<t_{1}$, the optimal algorithm on $G^{\prime}$ will not match any second-class vertex. So, we can force the optimal algorithm on $G$ to never match $u$ before time $t_{1}$. On the other hand, since the algorithm will decide to match a second-class online vertex after $t_{1}$, we can also force it to match to $u$ as long as possible. With these two constraints, the algorithm still achieve an expected objective of $\textsc{Opt}(G,t)$, and on $u$ it actually acts the same as the optimal algorithm on $G^{\prime}$, which means $u$ contributes exactly $\textsc{Opt}(G^{\prime},t)$ to the expected objective.
Now consider an algorithm for some $t^{\prime}<t$ which satisfies the two constraints and always match a first-class vertex if possible. It’s straight-forward to see that $u$ must contribute $\textsc{Opt}(G^{\prime},t^{\prime})$. On the other hand, at any time in $[t,1]$, the probability that $u$ has been matched in our algorithm is at least the probability that $u$ has been matched in the optimal algorithm for $t$. So, we can simply do not match $v$ in $[t^{\prime},t)$, and then in $[t,1]$ match $v$ in the same way as the optimal algorithm for $t$. Then the expected objective of our algorithm is $\textsc{Opt}(G^{\prime},t^{\prime})+\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$, which means $\textsc{Opt}(G,t^{\prime})-\textsc{Opt}(G^{\prime},t^{\prime})\geq\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$. In other words, $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ is decreasing in $[0,t_{1}]$ as desired. ∎
Combining 1 and 3, we can get our hardness result.
@@
For any $k,t_{0},t_{1}$, let $\textsc{Alg}(k,t_{0},t_{1})$ denote the expected objective of Algorithm 2. According to our matching strategy, $\textsc{Alg}(k,t_{0},t_{1})$ can be calculated using straightforward calculus. However, the closed-form expression is super long, so we omit it here. See Section A.1 for computational details.
By 1, the optimal LP value is $2\ln 2+(2-\ln 2)k$. So the competitive ratio is $\frac{\textsc{Alg}(k,t_{0},t_{1})}{2\ln 2+(2-\ln 2)k}$. For $k\approx 3.40216$, the ratio achieves its maximum value $\approx 0.66275$ at $t_{0}\approx 0.12437$ and $t_{1}\approx 0.29539$. ∎

### v2

@@
The online stochastic matching problem was introduced by [feldman2009online], together with the $(1-\frac{1}{e})$-competitive Suggested Matching algorithm. In the most general edge-weighted setting, this ratio has not been improved for more than one decade, until recently [yan2024edge] beat the $1-\frac{1}{e}$ bound and [qiu2023improved] further improved it to $0.650$. Both works measure the online competitiveness against the offline LP relaxation introduced by Jaillet and Lu [jaillet2014online]. The same LP has also played an important role in other settings as it is a natural choice for two-choice online algorithms.
In this paper, we prove an upper bound of $0.663$ and a lower bound of $0.662$ for edge-weighted online stochastic matching under Jaillet-Lu LP. We propose a simple hard instance and identify the optimal online algorithm for this specific instance which has a competitive ratio of <0.663<0.663. Despite the simplicity of the instance, we then show that a near-optimal algorithm for it, which has a competitive ratio of >0.662>0.662, can be generalized to work on all instances without any loss.
As our algorithm is generalized from a real near-optimal algorithm instead of manually combining trivial strategies, it has two natural advantages compared with previous works: (1) its matching strategy varies from time to time; (2) it utilizes global information about offline vertices. On the other hand, the upper bound suggests that more powerful LPs and multiple-choice strategies are needed if we want to further improve the ratio by >0.001>0.001.
@@ 1 Introduction
The online bipartite matching problem was introduced by [karp1990optimal], motivated by applications such as online advertising. In this problem, we are given a bipartite graph. Advertisers are modeled as offline vertices on one side of the graph, which are known at the beginning. Impressions (user queries) are modeled as online vertices on the other side, which arrive one by one. The weights on the edges indicate how much the advertisers are interested in the impressions. When an online vertex arrives, the online algorithm sees its incident edges, and must immediately and irrevocably decide how to match it. [karp1990optimal] proposed the celebrated Ranking algorithm, which achieves a competitive ratio of $1-\frac{1}{e}\approx 0.632$ in the unweighted setting. This is optimal in the worst case model, where we assume the online algorithm has no prior information about the graph.
@@
However, in real-world scenarios, the algorithm often knows some prior information. For example, online advertising platforms can use historic data to predict user behaviors in the future. [feldman2009online] introduced the online stochastic matching model, where we assume all online vertices are drawn independently from an identical distribution which is known by the algorithm in the beginning. [feldman2009online] proposed the Suggested Matching algorithm which is $(1-\frac{1}{e})$-competitive even in the most difficult edge-weighted setting, and then beat the $1-\frac{1}{e}$ bound in the unweighted setting under the integral assumption [^1] It means that the expected number of online vertices of each type is an integer.. Further works removed the integral assumption [manshadi2012online, jaillet2014online, huang2021online, huang2022power, tang2022fractional] and beat the $1-\frac{1}{e}$ bound in the vertex-weighted setting [huang2021online, huang2022power, tang2022fractional] and the free-disposal [^2] It allows us to match an offline vertex many times and only count the most valuable one in the end. edge-weighted setting [huang2022power].
@@
A common framework of the online algorithms is to solve an LP to get a fractional matching which upper-bounds the offline optimal solution, and then use the fractional solution to guide the online strategy. The online competitiveness is then measured against the LP relaxation. In particular, the LP proposed by Jaillet and Lu [jaillet2014online] is quite useful for two-choices algorithms (which most previous algorithms are) in the sense that it limits the case that an online vertex type is mostly matched to one offline vertex in the fractional solution, forcing most online vertices to have at least two choices. Under Jaillet-Lu LP, [huang2022power] provided both a lower bound and an upper bound of the same ratio 0.7060.706333The exact value is $1-\frac{1}{1-\ln 2}\big(\frac{1}{2e}-\frac{\ln 2}{e^{2}}\big)$. for unweighted, vertex-weighted and free-disposal edge-weighted settings.
@@
The edge-weighted setting was known to be strictly harder than the above settings [huang2022power]. Recently, [yan2024edge] beat the $1-\frac{1}{e}$ bound with a competitive ratio of $0.645$ by extending the Suggested Matching algorithm to have different strategies in $3$ different time periods. [qiu2023improved] improved the ratio to $0.650$ by increasing the number of periods to $5$. Both algorithms still use the trivial strategy from Suggested Matching in more than half of the time. In contrast, instead of using several fixed strategies in different periods, the state-of-the-art Poisson Online Correlated Selection algorithm [huang2022power] for unweighted/vertex-weighted settings has a strategy which smoothly changes over time. A natural question arises: Can smoothly changed strategies help to further improve the competitive ratio in the edge-weighted setting?
@@
On the other hand, both algorithms [yan2024edge, qiu2023improved] use the Jaillet-Lu LP relaxation. We know that the provably best ratio for online algorithms under Jaillet-Lu LP is $0.706$ in the other $3$ settings. What about the edge-weighted setting?
@@ 1.1 Our Results
We first propose a hard instance and show that, for this instance, the optimal online algorithm can only achieve a competitive ratio <0.663<0.663. Then, we extract some key properties of a near-optimal matching process of this specific instance which are sufficient to guarantee a competitive ratio >0.662>0.662. Finally, we propose a generalized algorithm which maintains these properties for all instances, which is therefore $0.662$-competitive for all instances. The algorithm improves the best known ratio of $0.650$ [qiu2023improved], and the hardness result suggests that more powerful LPs are necessary for further improvements. [^4] In the unweighed and vertex-weighted settings, [huang2022power] goes beyond the optimal algorithm against Jaillet-Lu LP by utilizing more powerful LPs. In the free-disposal edge-weighted setting, there is still no algorithm achieving this.
Compared with existing algorithms [yan2024edge, qiu2023improved] for edge-weighted online stochastic matching, our algorithm has two major differences: (1) instead of only using one or several fixed strategies, we deploy a strategy that smoothly varies from time to time; (2) when an online vertex arrives, instead of only looking at its neighbors locally, our algorithm utilizes some global information about the probability distribution of remaining offline vertices. Intuitively speaking, both are natural movements towards the optimal algorithm.
@@
Previous work [yan2024edge] showed that, under Jaillet-Lu LP, the problem can be reduced to the case that there are only two classes of online vertex types:
@@
In Section 3, we identify the optimal online algorithm for this instance. By carefully setting the edge weights, we show that the optimal algorithm achieves a competitive ratio <0.663<0.663.
@@
After getting the optimal algorithm for the specific hard instance, we try to design our algorithm by generalizing the special algorithm to work on all instances. For each first-class (resp. second-class) edge, we want to match it at the same rate as the first-class (resp. second-class) edges in the hard instance. First-class vertices are easy to deal with since they have only one possible choice when arriving, so we focus on second-class vertices. When a second-class vertex arrives, to make sure that it is feasible to match it at the desired rate, we (at least) need an upper bound of the probability that the two offline neighbors have both been matched.
To be able to simulate the algorithm on all instances, this probability must take its maximum value in our hard instance, so that the upper bound works for all instances. In other words, roughly speaking, the algorithm should match two offline vertices in the hard instance in the most positively correlated way. Unfortunately, the optimal algorithm for the hard instance does not have this property. Therefore, in Section 4, we introduce a near-optimal algorithm with this property which has a competitive ratio >0.662>0.662 on the hard instance. Then we show that the near-optimal algorithm can be generalized to have the same competitive ratio on all instances.
@@ 1.2 Other Related Work
In the worst case model, the Ranking algorithm [karp1990optimal] and its generalization [aggarwal2011online] achieved the optimal competitive ratio of $1-\frac{1}{e}$ for unweighted and vertex-weighted matching. For free-disposal edge-weighted matching, a sequence of works [fahrbach2020edge, gao2022improved, blanc2022multiway] improved the competitive ratio to $0.5239$.
In the random order model [^5] The adversary decides the graph, then online vertices arrive in a uniformly random order., the currently best ratios are $0.696$ [goel2008online, mahdian2011online] for unweighted matching and $0.662$ [huang2019online, jin2021improved] for vertex-weighted matching.
In the online stochastic matching model, for unweighted and vertex-weighted matching, a sequence of works [feldman2009online, bahmani2010improved, manshadi2012online, jaillet2014online, huang2021online, huang2022power, tang2022fractional] improved the competitive ratio to $0.716$ and the upper bound to $0.823$. For free-disposal edge-weighted matching, [huang2022power] proposed a $0.706$-competitive algorithm. Under the integral assumption, the state-of-the-art algorithms achieved competitive ratios of $0.7299$ for unweighted matching [manshadi2012online, jaillet2014online, brubach2020online], $0.725$ for vertex-weighted matching [jaillet2014online], and $0.705$ for edge-weighted matching [haeupler2011online, brubach2020online]. An upper bound of $0.703$ [huang2022power] for (general) edge-weighted matching inferred that it is strictly harder than the above settings. For non i.i.d. arrivals, recent works [tang2022fractional, chen2024stochastic] achieved a competitive ratio of $0.69$ for unweighted and vertex-weighted matching.
@@ 2 Preliminaries
In the Poisson arrival model, instead of assuming a deterministic number of online vertices, we assume that online vertices arrive following a Poisson process with arrival rate $\Lambda$ in the time horizon $[0,1]$. An equivalent view is that each online vertex type $i\in I$ independently follows a Poisson process with arrival rate $\lambda_{i}$. This equivalency and the continuous arrival process greatly help the analyses of online stochastic matching algorithms, after [huang2021online] proved that the Poisson arrival model is asymptotically equivalent to the original model. In this paper, we present our upper and lower bounds in the Poisson arrival model. We remark that the same bounds hold in the original model with an additive term of $o(1)$ with respect to $\Lambda$. We also remark that prior works in the original model also conventionally assume a large enough $\Lambda$, or explicitly have the $o(1)$ additive term.
@@
The (Poisson arrival version of) Jaillet-Lu LP [jaillet2014online] is as follows:
@@
A solution of the Jaillet-Lu LP is a fractional matching between online vertex types and offline vertices. [huang2021online] proved that the optimal value of the LP, that is, the objective of the optimal fractional matching, upper-bounds the expected objective of the optimal offline integral matching. In fact, each $x_{ij}$ upper-bounds the probability that offline vertex $j$ is matched to an online vertex of type $i$. The reason behind Equation 1 is that, an online vertex type cannot put too much of its arrival rate to the same offline vertex since the probability that it arrives (at least once) is smaller than the arrival rate.
In this paper, we consider the competitive ratio against the optimal value of the Jaillet-Lu LP instead of the optimal integral matching.
Without loss of generality, we define $x_{ij}=0$ for any pair $(i,j)\notin E$. For each online vertex type $i$, define $x_{i}=\sum_{j\in J}x_{ij}$. For each offline vertex $j$, define $x_{j}=\sum_{i\in I}x_{ij}$.
@@
[yan2024edge] showed that, any instance $A$ can be efficiently transformed into another instance $B$ such that the optimal fractional matching (under Jaillet-Lu LP) of $B$ satisfies the following conditions:
@@
In addition, for any online algorithm which works for $B$, using the transformation as a preprocessing step, it works for $A$ with the same competitive ratio. So, without loss of generality, in this paper we restrict the instance to satisfy the above conditions.
@@ 3 Hardness
The optimal value of Jaillet-Lu LP on $G$ is $2\ln 2+(2-2\ln 2)k$.
Now we are going to find the optimal online algorithm on $G$. Clearly, since $k\geq 1$, when a first-class online vertex arrives, we shall always match it if we can. What about a second-class online vertex? Intuitively, the decision should depend on $k$ and the current time $t$. When $t$ is small but $k$ is large, we may want to discard this online vertex even if it has unmatched neighbors, to preserve the chance of matching a second-class vertex in the future. Also notice that the number of unmatched neighbors may also affect our decision, since it affects the chance that the offline vertex can be matched with another second-class online vertex in the future.
@@
When a second-class online vertex arrives, the algorithm can match it if $\textsc{Opt}(G^{\prime},t)+1\geq\textsc{Opt}(G,k)$, and it doesn’t matter which neighbor is matched by symmetry. When $t>t_{1}$, we know $\textsc{Opt}(G^{\prime},t)\leq 1$, then $\textsc{Opt}(G^{\prime},t)+1\geq 2\textsc{Opt}(G^{\prime},t)\geq\textsc{Opt}(G,t)$. So the algorithm can always match it when $t>t_{1}$. Then, it remains to show that $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ is decreasing in $0,t_{1}]$. If this is true, then we can set $t_{0}$ as the boundary where $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ starts going below $1$.
@@
Note that, by Lemma 2, when $t<t_{1}$, the optimal algorithm on $G^{\prime}$ will not match any second-class vertex. So, we can force the optimal algorithm on $G$ to never match $u$ before time $t_{1}$. On the other hand, since the algorithm will decide to match a second-class online vertex after $t_{1}$, we can also force it to match to $u$ as long as possible. With these two constraints, the algorithm still achieves an expected objective of $\textsc{Opt}(G,t)$, and on $u$ it actually acts the same as the optimal algorithm on $G^{\prime}$, which means $u$ contributes exactly $\textsc{Opt}(G^{\prime},t)$ to the expected objective.
Now consider an algorithm for some $t^{\prime}<t$ which satisfies the two constraints and always match a first-class vertex if possible. It’s straight-forward to see that $u$ must contribute $\textsc{Opt}(G^{\prime},t^{\prime})$. On the other hand, at any time in $[t,1]$, the probability that $u$ has been matched in our algorithm is at least the probability that $u$ has been matched in the optimal algorithm for $t$. So, we can simply do not match $v$ in $t^{\prime},t)$, and then in $[t,1]$ match $v$ in the same way as the optimal algorithm for $t$. Then the expected objective of our algorithm is $\textsc{Opt}(G^{\prime},t^{\prime})+\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$, which means $\textsc{Opt}(G,t^{\prime})-\textsc{Opt}(G^{\prime},t^{\prime})\geq\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$. In other words, $\textsc{Opt}(G,t)-\textsc{Opt}(G^{\prime},t)$ is decreasing in $0,t_{1}]$ as desired. ∎
Combining 1 and 3, we get our hardness result.
@@
For any $k,t_{0},t_{1}$, let $\textsc{Alg}(k,t_{0},t_{1})$ denote the expected objective of Algorithm 2. According to our matching strategy, $\textsc{Alg}(k,t_{0},t_{1})$ can be calculated using straightforward calculus. However, the closed-form expression is extremely long, so we omit it here. See Section A.1 for computational details.
By 1, the optimal LP value is $2\ln 2+(2-2\ln 2)k$. So the competitive ratio is $\frac{\textsc{Alg}(k,t_{0},t_{1})}{2\ln 2+(2-2\ln 2)k}$. For $k\approx 3.40216$, the ratio achieves its maximum value $\approx 0.66275$ at $t_{0}\approx 0.12437$ and $t_{1}\approx 0.29539$. ∎